x+y=201
x-y=11 solve using substitution

Answer by lorst(1)

x+y=201
x-y=11
let:x-y=11
x=11+y
solve for y:
x+y=201
11+y+y=201
2y=201-11
2y/2=190/2
y=95
solve for x:
x-y=11
x-95=11
x=11+95
x=106
check:
x+y=201
106+95=201
201=201 check

x-y=11
106-95=11
11=11 check